OpenStudy (anonymous):

A boat makes a 120-mile trip downstream in 3 hours but makes the return trip in 4 hours. What is the rate of the current? 3 mph 4 mph 5 mph I need help setting up the problem

OpenStudy (anas.p):

First we take out the speed of both upstream and downstream

OpenStudy (anas.p):

Speed = Distance/Time So Downstream the speed is 120/3=40 40mph

OpenStudy (anas.p):

and upstream is 120/4

OpenStudy (anas.p):

=30

OpenStudy (anonymous):

which is 30... but how do i find the current speed though?

OpenStudy (anas.p):

the speed of the current will be 40-30=10

OpenStudy (anonymous):

ooh i see now, thanks

hartnn (hartnn):

but i am getting speed of boat as 35 mph and speed of current as 5 mph....:\

hartnn (hartnn):

@Anas.P ,u got speed of (boat + current) as 40 mph and speed of (boat - current)=30 mph.....isn't it? which makes current speed as 5mph....

OpenStudy (anonymous):

how does it make 5?

hartnn (hartnn):

too easy to use a calcy.... downstream means current will assist boat so total speed = their sum up stream means current will resist boat so total speed = their difference as distance covered is same we can equate product of time and speed 3*(boat+current) = 4* (boat - current) = 120 which when easily solved gives, boat = 35mph , current = 5 mph

hartnn (hartnn):

u got this or should i use variables x and y and solve again with equations?? @willow

OpenStudy (anas.p):

@hartnn thanks a lot, sorry i missed it. But can you also use variables to make it more clear..

OpenStudy (anonymous):

i think i understand, but can you use vairables?

hartnn (hartnn):

sure, let x be the speed of boat,let y be the speed of current Downstream equation (speed *time = dist ) (x+y)*3=120 (x+y)=40 ---> Equation 1 upstream equation (speed *time = dist ) (x-y)*4=120 (x-y)=30 ---> Equation 2 now from eq 2,x=30+y this x i will put in eq 1 so 30+y+y=40 so 2y=40-30 so y=10/2 so speed of current = 5mph okay? :)

OpenStudy (anas.p):

Cool thanks..